OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . .

OpenStudy (chihiroasleaf):

this is arithmetic sequence, the sum of the n first term is \[Sn = \frac{ n }{ 2 }\left( 2a + \left( n-1 \right) d \right)\] with n = the number of terms a = the first term d = the difference you can use this formula to solve the problem can you determine, what is n, a, and d of your sequence?

OpenStudy (anonymous):

d=-9-(-4) =-9+4 =-5 n=12 a=1.

OpenStudy (chihiroasleaf):

yess... :) now, plug in the value into the formula...

OpenStudy (anonymous):

12/2(2*1+(12-1)-5)

OpenStudy (anonymous):

which equals 48?

OpenStudy (chihiroasleaf):

be careful in calculation.... :)